10 chemists and 15 biologists apply for a position in a company. If 5 people are hired at random, what is the probability that 2 are chemists and 3 are biologists?
Answer by stanbon(75887)

10 chemists and 15 biologists apply for a position in a company. If 5 people are hired at random, what is the probability that 2 are chemists and 3 are biologists?
-----
# of ways to succeed: 10C2*15C3 = 45*455
---
# of random sets of 5: 25C5 = 53130
----
Ans: (45*455)/53130 = 0.3854
==================
Cheers,
Stan H.
==================